OpenStudy (anonymous):

Walter invested $15,000 in two accounts. The Rowe's Best account paid 6% while the Goings Great account paid 7%. If he earned $980 in interest for the year, how much did he invest in the Rowe's Best account?

OpenStudy (anonymous):

980/0.06

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

after I divide do I need to find the answer of .07? I got really confused

OpenStudy (anonymous):

I have different kind of answer everytime

OpenStudy (anonymous):

X=money invested in the Rowe's Best Account Y=money invested in the Goings Great account X+Y=15000 ->Y=15000-X X(6%)+Y(7%)=980 (0.06X)+(15000-X)(0.07)=980 Solve for X

OpenStudy (anonymous):

Does this make sense to you?

OpenStudy (anonymous):

the first solution I did I have 7500 now I got 6500.I dont know where did I do wrong here

OpenStudy (anonymous):

You can try to plug X and Y into X(6%)+Y(7%)=980 see if they are right

OpenStudy (anonymous):

(0.06X)+(15000-X)(0.07)=980 X should be = 7000 if u did the math right

OpenStudy (anonymous):

how did you get the 7000?can u show me the solution?

OpenStudy (anonymous):

0.06X+1050-0.07X=980 -0.01X=-70 X=7000

OpenStudy (anonymous):

okey thats help me how you do it thank you so much

OpenStudy (anonymous):

No Problem :)